simplify 2z^3 -6z^2 over z^2 - 3z. identify any z-values for which the expression is undefined.
1
Expert's answer
2012-04-17T08:14:45-0400
(2z^3 -6z^2)/(z^2 - 3z)=z^2(2z -6)/{z(z - 3)}=2z^2(z -3)/{z(z - 3)}=2z defined for any real z